# Lyman Continuum leakage from massive leaky starbursts: A different class of emitters?

### v1

@@ 3.3 Measurement of ancillary parameters
This paper aims to study the LyC escape fraction and its dependence on host galaxy properties compared with the LzLCS survey. Here we list the ancillary parameters we measure.
We measure the equivalent width (EW) of the Ly emission line from the extracted COS spectra. The continuum is modeled by fitting a linear function to the continuum within a spectral window of 30Å on either side of the emission line using an iterative sigma clipping technique to exclude noise spikes. To measure Ly EW, the observed spectrum is normalized by dividing the spectrum by the modeled continuum and then corrected for redshift. All measured fluxes and EWs of the rest-frame optical emission lines are taken from the Sloan Digital Sky Survey (SDSS) spectral data, specifically from the value-added catalog provided by the Portsmouth group (Thomas et al., 2013). A signal-to-noise ratio (S/N) cut of 5 is imposed for any emission line detection. The flux values for [OIII]5007 and [OII]3726, 3729Å emission lines are obtained from the catalog to compute the [OIII]/[OII] ratio.
These are dusty systems, so our primary measure of SFR is from the WISE mid-IR photometry, as described above. We also calculated based on the observed continuum flux measured at 1500Å (see above). We do not use a SFR based on the Balmer emission-lines since these can significantly underestimate the SFR in cases in which only a fraction of the ionizing photons are absorbed by HI in the galaxy ISM.
@@
where . The relation is valid for -1 O3N3 1.9 and is insensitive to dust extinction. Note, for the LzLCS+ sample, the oxygen abundance is calculated form the [OIII]4363Å line.
@@ 4.1 LyC escape fraction
We detect a significant flux in the LyC region below the Lyman edge for 3 out of our 5 galaxies. For these three detected sources, we report the mean background-subtracted flux density in the LyC spectral window (as explained in §3.2). In case of non-detections, we take the 84th percentile of the background distribution to be the upper limit on following Flury et al. (2022a). To confirm the LyC detections, we examine the cross-dispersion profiles of the two-dimensional spectra of our sources in the LyC window and compare with the non-ionizing stellar counterpart centered at 1100 Å. Fig. 4 shows example profiles for a LyC detected and non-detected source.
For the LyC detected sources, we report flux densities () of 5.38 Å-1 for J1145, 2.9 Å-1 for J0121, and 1.58 Å-1 for J1230. For the other two sources, we report upper limits. The values are listed in Table. 1. Fig. 5 shows the observed LyC flux vs. redshift for the massive leakers sample (see §2 for the sample definition). Yellow square denote sources from this work, while yellow diamond represent Wang et al. (2019) and Borthakur et al. (2014) objects, compared with LzLCS+ objects in blue circles (Flury et al., 2022b).
As mentioned in §3.2, we report three different ways to measure the escape fraction. The first and the simplest method is the ratio of the observed fluxes in the LyC region to that of rest wavelength of 1100 Å /Fλ1100. The rest-frame 1100 Å flux provides a measurement for the non-ionizing stellar continuum, and thus, the mean ratio of LyC to 1100 Å is an indirect proxy for . It is very useful since it is directly derived from observed spectrum in a model-independent way without any assumptions about stellar populations or other galaxy properties.
@@ 4.2 Comparison with other indirect diagnostics
The LzLCS survey and other studies in the past have shown that the LyC escape often exhibits significant correlation with different host galaxy properties, albeit with substantial scatter (Wang et al., 2021; Flury et al., 2022b; Izotov et al., 2016b; Saldana-Lopez et al., 2022; Chisholm et al., 2022; Jaskot et al., 2024a). The observed correlations with , e.g. with , , Ly EW, H EW, [OIII]/[OII] (), dust attenuation, [SII]-deficiency, -slope and residual flux in Ly motivated some of these quantities to be used as potential indirect signposts for leakage. These indirect tracers are especially important for sources where directly measuring LyC flux density is impossible, like during the EoR (e.g., Jaskot et al., 2024b). Here, we evaluate these correlations, which have been extensively explored in traditional LyC-leaking galaxies from the LzLCS+ sample, to determine their relevance to the massive leaker population. For this comparison, we use two different proxies for LyC escape in the Y-axis. Fig. 8 shows the direct flux ratio (or ) as it is uniformly reported in previous studies and is independent of the specific method used to calculate . Fig. 9 shows the variation of those same quantities but with LyC escape fraction derived from stellar population modeling, which represents the total fraction of LyC photons escaping the IGM after accounting for both neutral hydrogen absorption and dust attenuation.
Fig 8 illustrates the correlations between and these different indirect tracers of LyC escape. Each panel displays the Kendall coefficient, indicating the strength of the correlation between the x and y values, along with the associated p-value representing the probability of the null hypothesis. Black-colored text represent the Kendall coefficients for the LzLCS+ sample alone, while red-colored values include both the LzLCS+ sample and the massive leakers. The results clearly show that the significance of the correlations (as indicated by ) decreases when massive leakers are included (red) compared to the LzLCS+ sample alone (black). Visually, it is also evident that the massive leakers exhibit substantial deviations from the LzLCS+ leakers across a broad range of parameters.
@@
The upper left panel shows the relation between and the Ly equivalent width (EW) for the LzLCS+ sources (blue circles) and the massive leakers (yellow diamonds and squares). There is a strong positive correlation between the two quantities for the LzLCS+ sources. The Ly line is produced from resonant scattering of Ly photons through the line of sight, and it is plausible for it to follow , since both Ly and LyC emission depend on the distribution of neutral hydrogen in a galaxy. Hence Ly EW is often used as an indirect estimate (e.g., Roy et al., 2023). However, the main result here is that the massive leaker sample is shifted towards lower Ly EW compared to the LzLCS+ galaxies. The mean Ly EW value for the former is 17 Å, which is 4 times lower compared with the mean Ly EW = 70 Å from the LzLCS+ sample. The lower middle panel is an analogous plot for the EW. Here the massive leakers are even more strongly offset from the LzLCS+ galaxies, with much weaker H emission.
@@
Fig. 8 (lower left panel) show a rather tight correlation between the residual flux (RF) in the core of the absorption-line and /Fλ1100. Like [SII], this correlation includes both type of leakers. The mean Ly RF for the massive leakers sample = 0.346, which is roughly two times higher than that of the LzLCS+ sample (Saldana-Lopez et al., 2022). The residual fluxes (RF) at the bottom of the HI absorption lines are tracers of the covering fraction of the neutral hydrogen (). In the case of a uniform HI screen geometry, the residual flux RF = 1 . As illustrated by Heckman et al. (2011), this observed trend is also consistent with picket-fence scenario where the covering factor of optically thick HI is not unity. Supernova-driven outflows or extremely intense LyC radiation (large U) can create channels or “holes” in the HI through which ionizing radiation can escape, leading to higher .
@@
Fig. 9 presents the same correlations as Fig. 8, but with the LyC escape fraction , derived from stellar population modeling (Method 2 in §3.2), replacing . The changes observed on the Y-axis compared to Fig. 8 reflect the significant impact of dust attenuation on the resulting escape fraction that ultimately escapes into the IGM, as highlighted previously.
@@ 4.3 Are these a different class of leakers?
We agree with the hypothesis put forward by Heckman et al. (2011); Borthakur et al. (2014); Alexandroff et al. (2015); Wang et al. (2019); Flury et al. (2022b); Jaskot et al. (2024a) that the LyC escape in these massive leakers is made possible by the extreme feedback effects produced by an extremely massive and compact starburst (Flury et al., 2024). In the three massive leakers discussed by Borthakur et al. (2014) and Wang et al. (2019), the UV light from starburst was dominated by highly compact (marginally resolved by HST), massive objects located at or near the galactic nucleus. These were dubbed Dominant Central Objects (DCOs; Overzier et al., 2009). In this study, we have confirmed LyC emission in three out of the five sources. All three have very high SFR/Area, but only J0121 has the morphology of a compact DCO as shown in Fig. 1. The rest have comparatively extended morphology. This suggest that the more general property of very high SFR/Area is responsible for enabling the escape of LyC photons. The typically large values of the RF in the absorption-line in the massive leakers is fully consistent with a “picket fence” geometry for the HI with holes that are fully ionized, enabling LyC escape and also reducing the relative amount of partially-ionized gas traced by [SII]. It is noteworthy that high SFR/A is correlated with high velocities in the starburst-generated galactic winds (e.g., Heckman et al., 2015, 2017; Davis et al., 2023). This suggests that the ionized holes may be created - at least in part - as a result of the extremely high ram pressure associated with the wind fluid. Indeed, Amorín et al. (2024) recently found correlations between the velocity of these feedback driven winds and the LyC escape fraction, while Flury et al. (2024) explored the in-depth connections between the different types of feedback mechanism and the LyC escape. Our low-resolution HST/COS data could not confirm blueshifts in the absorption lines indicative of outflows. We are currently planning follow-up observations with the Large Binocular Telescope to obtain higher spectral resolution data. This analysis will be addressed in future work.
@@ 5 Conclusions
We present HST/COS spectroscopy observations of five low-redshift (z 0.3) starburst galaxies, selected for their high stellar mass () and relatively weak [S II] 6717Å,6731Å nebular emission lines compared to typical star-forming galaxies. The weakness in [S II] is hypothesized to indicate galaxies that are optically thin to ionizing radiation. Recently, several high-redshift galaxies (z > 5) observed with JWST have also shown extremely weak [S II] emission (Sanders et al., 2023; Cameron et al., 2023), reinforcing the potential of weak [S II] as a marker of LyC escape in high-z galaxies.
Significant Lyman continuum (LyC) flux was detected in three of the five galaxies, with dust-corrected escape fractions () of 71%, 84%, and 33%, representing the fraction of LyC photons that would escape in the absence of dust. However, the total escape fraction, which is the real fraction of photons escaping into the IGM the effects of both hydrogen absorption and dust extinction, is significantly lower, ranging from 1-3%.
We compared these [S II]-weak galaxies to other known classes of LyC-leaking galaxies. In contrast to the low-redshift traditional leakers from the LzLCS+ sample, these galaxies exhibit higher stellar masses, greater metallicities, more dust extinction, lower ionization states (as indicated by [OIII]/[OII]), smaller Ly and H equivalent widths, and higher star formation rate surface densities. These results indicate that the massive leakers in our sample constitute a distinct population from the traditional LzLCS+ leakers, suggesting the presence of multiple mechanisms facilitating LyC photon escape. The characteristics of these massive leakers align with a “picket-fence” model, in which intense feedback creates channels or holes in the neutral hydrogen, allowing LyC photons to escape through these openings. This expands the range of galaxy properties over which searches for LyC-leaking galaxies can be conducted, enhancing our ability to use low-redshift leakers as local analogs for studying the physical processes that allow LyC photons to escape. It also points to the likelihood that various physical conditions and mechanisms contribute to LyC leakage.

### v2

@@ 3.3 Measurement of ancillary parameters
This paper aims to study the LyC escape fraction and its dependence on host galaxy properties compared with the LzLCS+ survey. Here we list the ancillary parameters we measure.
We measure the equivalent width (EW) of the Ly emission line from the extracted COS spectra. The continuum is modeled by fitting a linear function to the continuum within a spectral window of 30Å on either side of the emission line using an iterative sigma clipping technique to exclude noise spikes. To measure Ly EW, the observed spectrum is normalized by dividing the spectrum by the modeled continuum and then corrected for redshift. The fluxes and EWs of the rest-optical emission lines are taken from Sloan Digital Sky Survey (SDSS) spectra, as listed in the value-added catalog produced by the Portsmouth group (Thomas et al., 2013). A signal-to-noise ratio (S/N) cut of 5 is imposed for any emission line detection. The flux values for [OIII]5007 and [OII]3726, 3729Å emission lines are obtained from the catalog to compute the [OIII]/[OII] ratio.
These are dusty systems, so our primary measure of SFR is based on both from the WISE mid-IR photometry, and based on the observed continuum flux measured at 1500Å (see above). We do not use a SFR based on the Balmer emission-lines since these can significantly underestimate the SFR in cases in which only a fraction of the ionizing photons are absorbed by HI in the galaxy ISM.
@@
where . The relation is valid for -1 O3N3 1.9 and is insensitive to dust extinction. Note, for the LzLCS+ sample, the oxygen abundance is calculated form the [OIII]4363Å line. We provide these measurements of the host galaxy properties of our sources in Table. 2.
@@ 4.1 LyC escape fraction
We detect a significant () flux in the LyC region below the Lyman edge for 3 out of our 5 galaxies. For these three detected sources, we report the mean background-subtracted flux density in the LyC spectral window (as explained in §3.2). In case of the two non-detections, we take the 84th percentile of the background distribution to be the upper limit on following previous LzLCS+ studies (Flury et al., 2022a). The corresponding uncertainties in the reported LyC fluxes are derived by summing the two error sources in quadrature: the errors from the HST/COS spectra within the LyC spectral window and the uncertainties from bootstrapped residual dark current fluxes across the entire detector region. To confirm the LyC detections, we also examine the cross-dispersion profiles of the two-dimensional spectra of our sources in the LyC window and compare with the dark current flux levels and the non-ionizing stellar counterpart centered at 1100 Å. Fig. 4 shows example profiles for a LyC detected and non-detected source.
For the LyC detected sources, we report flux densities (and errors) of 5.38 ( Å-1 for J1145, 2.9 ( Å-1 for J0121, and 1.58 Å-1 for J1230. The detection significances for the three LyC-detected sources, computed using FeldCous (see §3.2), are listed in Table 1. For the two non-detected sources that fall below the threshold, we report upper limits. Following the classification scheme of Flury et al. (2022a), two of our LyC leakers (J0121 and J1230) are classified as weak detections (), while J1145 qualifies as a strong LyC source with significance. In Fig. 5, we present the LyC fluxes as a function of redshift for our massive leaker sample (yellow squares), alongside previously published data on similarly massive LyC leakers from Wang et al. (2019) and Borthakur et al. (2014) (yellow diamonds). These are compared to the traditional LzLCS+ sample, as detailed in Flury et al. (2022b) (see §2 for a comprehensive description of the sample).
As mentioned in §3.2, we report three different ways to measure the escape fraction. The first and the simplest method is computing the ratio of the observed fluxes in the LyC region to that of rest wavelength of 1100 Å /Fλ1100. We denote this as method 1. The rest-frame 1100 Å flux provides a measurement for the non-ionizing stellar continuum, and thus, the mean ratio of LyC to 1100 Å is an indirect proxy for . It is very useful since it is directly derived from observed spectrum in a model-independent way without any assumptions about stellar populations or other galaxy properties.
@@ 4.2 Comparison with other indirect diagnostics
The LzLCS+ survey and other studies in the past have shown that the LyC escape often exhibits significant correlation with different host galaxy properties, albeit with substantial scatter (Wang et al., 2021; Flury et al., 2022b; Izotov et al., 2016b; Saldana-Lopez et al., 2022; Chisholm et al., 2022; Jaskot et al., 2024a). The observed correlations with , e.g. with , , Ly EW, H EW, [OIII]/[OII] (), dust attenuation, [SII]-deficiency, -slope and residual flux in Ly motivated some of these quantities to be used as potential indirect signposts for leakage. These indirect tracers are especially important for high redshift sources where directly measuring LyC flux density is impossible, for example, during the EoR (e.g., Jaskot et al., 2024b). In this analysis, we assess whether these correlations, previously established in traditional LyC-leaking galaxies from the LzLCS+ sample, also apply to the massive leaker population. For this comparison, we use three distinct proxies for LyC escape on the Y-axis.
Fig. 8 displays the relationships between these indirect tracers and the LyC flux ratio (or ), a model-independent, directly observed metric that is not directly affected by dust but does not directly measure . Fig. 9 presents these tracers in relation to the more traditional LyC escape fraction , derived through stellar population modeling. This widely used measure of accounts for attenuation effect of LyC photons from both neutral hydrogen absorption and dust. Finally, Fig. 10 shows the variation with , which isolates the effects of HI absorption alone. Each panel in these three figures displays the Kendall coefficient, indicating the strength of the correlation between the x and y values, along with the associated p-value representing the probability of the null hypothesis. Black-colored text represent the Kendall coefficients for the LzLCS+ sample alone (blue circles), while red-colored values include both the LzLCS+ sample and the massive leakers (yellow diamonds: literature sample from Wang et al., 2019; Borthakur et al., 2014, yellow squares: this work). The difference in the values show whether the significance of the correlations decreases when massive leakers are included compared to the LzLCS+ sample alone. This, in turn, quantifies whether the massive leakers lie along the main relation shown by the LzLCS+, or whether they are significantly offset.
@@
The upper left panel in Fig 8 shows the relation between and the Ly equivalent width (EW) for the LzLCS+ sources (blue circles) and the massive leakers (yellow diamonds and squares). There is a strong positive correlation between the two quantities for the LzLCS+ sources. The Ly line is produced from resonant scattering of Ly photons through the line of sight, and it is plausible for it to follow , since both Ly and LyC emission depend on the distribution of neutral hydrogen in a galaxy. Hence Ly EW is often used as an indirect estimate (e.g., Roy et al., 2023). However, the main result here is that the massive leaker sample is shifted towards lower Ly EW compared to the LzLCS+ galaxies. The mean Ly EW value for the former is 17 Å, which is 4 times lower compared with the mean Ly EW = 70 Å from the LzLCS+ sample. The lower middle panel is an analogous plot for the EW. Here the massive leakers are even more strongly offset from the LzLCS+ galaxies, with much weaker H emission.
@@
Fig. 8 (lower left panel) show a rather tight correlation between the residual flux (RF) in the core of the absorption-line and /Fλ1100. Like [SII], this correlation is maintained in both type of leakers. The mean Ly RF for the massive leakers sample = 0.346, which is roughly two times higher than that of the LzLCS+ sample (Saldana-Lopez et al., 2022). The residual fluxes (RF) at the bottom of the HI absorption lines are tracers of the covering fraction of the neutral hydrogen (). In the case of a uniform HI screen geometry, the residual flux RF = 1 . As illustrated by Heckman et al. (2011), this observed trend is also consistent with picket-fence scenario where the covering factor of optically thick HI is not unity. Supernova-driven outflows or extremely intense LyC radiation (large U) can create channels or “holes” in the HI through which ionizing radiation can escape, leading to higher .
@@
In summary, the massive leakers align with the LzLCS+ trend and their measurements only in the parameter spaces of vs. [SII]-weakness and Ly residual flux. Thus the massive leakers lie along the main relation for these two indirect tracers, but show a distinct offset in all the remaining quantities. A similar pattern appears in Fig. 10, which compares these tracers with instead of . Both these quantities remove or minimize the effects of dust attenuation. Thus, the massive leaker population have continuity with the traditional leakers in the two parameters most related to the effects of LyC absorption by HI: [SII]-weakness and Ly residual flux. Fig. 9, on the other hand, replaces with in the vertical axis, which is sensitive to both dust and HI. This comparison reveals that all quantities for the massive leakers distinctly deviate from the LzLCS+ sources, indicating that dust attenuation shifts these massive leakers away from the earlier trends observed in traditional leakers.
@@ 4.3 Are these a different class of leakers?
We agree with the hypothesis put forward by Heckman et al. (2011); Borthakur et al. (2014); Alexandroff et al. (2015); Wang et al. (2019); Flury et al. (2022b); Jaskot et al. (2024a) that the LyC escape in these massive leakers is made possible by the extreme feedback effects produced by an extremely massive and compact starburst (Flury et al., 2024). In the three massive leakers discussed by Borthakur et al. (2014) and Wang et al. (2019), the UV light from starburst was dominated by highly compact (marginally resolved by HST), massive objects located at or near the galactic nucleus. These were dubbed Dominant Central Objects (DCOs; Overzier et al., 2009). In this study, we have confirmed LyC emission in three out of the five sources. All three have very high SFR/Area, but only J0121 has the morphology of a compact DCO as shown in Fig. 1. The rest have comparatively extended morphology. This suggest that the more general property of very high SFR/Area is responsible for enabling the escape of LyC photons. The typically large values of the RF in the absorption-line in the massive leakers is fully consistent with a “picket fence” geometry for the HI with holes that are fully ionized, enabling LyC escape and also reducing the relative amount of partially-ionized gas traced by [SII]. It is noteworthy that high SFR/A is correlated with high velocities in the starburst-generated galactic winds (e.g., Heckman et al., 2015, 2017; Davis et al., 2023). This suggests that the ionized holes may be created - at least in part - as a result of the extremely high ram pressure associated with the wind fluid. Indeed, Amor´ın et al. (2024) recently found correlations between the velocity of these feedback driven winds and the LyC escape fraction, while Flury et al. (2024) explored the in-depth connections between the different types of feedback mechanism and the LyC escape. Our low-resolution HST/COS data could not confirm blueshifts in the absorption lines indicative of outflows. We are currently planning follow-up observations with the Large Binocular Telescope to obtain higher spectral resolution data. This analysis will be addressed in future work.
@@ 5 Conclusions
We present HST/COS spectroscopy observations of five low-redshift (z 0.3) starburst galaxies, selected for their high stellar mass () and relatively weak [S II] 6717Å,6731Å nebular emission lines compared to typical star-forming galaxies. The weakness in [S II] is hypothesized to indicate galaxies that are optically thin to ionizing radiation. Recently, several high-redshift galaxies (z 5) observed with JWST have also shown extremely weak [S II] emission (Sanders et al., 2023; Cameron et al., 2023), reinforcing the potential of weak [S II] as a marker of LyC escape in high-z galaxies.
Significant LyC flux was detected in three of the five galaxies, with dust-corrected escape fractions () of 71%, 84%, and 33%, indicating the proportion of LyC photons that would escape in dust-free conditions. However, dust substantially reduces the escaping LyC photon count. Consequently, the total escape fraction ()—the fraction of photons that actually escape and reach the IGM after surviving both neutral hydrogen absorption and dust extinction—is considerably lower, ranging from 1-3%.
We compared these [S II]-weak galaxies to other known classes of LyC-leaking galaxies. In contrast to the low-redshift traditional leakers from the LzLCS+ sample, these galaxies exhibit higher stellar masses, greater metallicities, more dust extinction, lower ionization states (as indicated by [OIII]/[OII]), smaller Ly and H equivalent widths, and higher star formation rate surface densities. These results indicate that the massive leakers in our sample constitute a distinct population from the traditional LzLCS+ leakers, suggesting the presence of multiple mechanisms facilitating LyC photon escape. The characteristics of these massive leakers align with a “picket-fence” model, in which intense feedback creates channels or holes in the neutral hydrogen, allowing LyC photons to escape through these openings. With JWST and other recent ground-based studies uncovering an increasing number of massive () galaxies at high redshifts () (Labbé et al., 2023; Long et al., 2024; Carnall et al., 2023; Nanayakkara et al., 2024; Carnall et al., 2020; Schreiber et al., 2018; Glazebrook et al., 2017), this population of massive leaky starbursts will expand the range of galaxy properties available for LyC-leakage studies. This also enhances our ability to use low-redshift leakers as local analogs for investigating the physical mechanisms that facilitate LyC photon escape at higher redshifts. It points to the likelihood that various physical conditions and mechanisms contribute to LyC leakage.